# Threads of Subtlety: Detecting Machine-Generated Texts Through Discourse Motifs

### v1

@@ Abstract
With the advent of large language models (LLM), the line between human-crafted and machine-generated texts has become increasingly blurred. This paper delves into the inquiry of identifying discernible and unique linguistic properties in texts that were written by humans, particularly uncovering the underlying discourse structures of texts beyond their surface structures. Introducing a novel methodology, we leverage hierarchical parse trees and recursive hypergraphs to unveil distinctive discourse patterns in texts produced by both LLMs and humans. Empirical findings demonstrate that, although both LLMs and humans generate distinct discourse patterns influenced by specific domains, human-written texts exhibit more structural variability, reflecting the nuanced nature of human writing in different domains. Notably, incorporating hierarchical discourse features enhances binary classifiers’ overall performance in distinguishing between human-written and machine-generated texts, even on out-of-distribution and paraphrased samples. This underscores the significance of incorporating hierarchical discourse features in the analysis of text patterns. The code and dataset will be available at [TBA].
@@
[…]
@@ 1 Introduction
On the contrary, LLMs simply generate texts sequentially based on the next token prediction task where the most likely next token in a sequence is predicted given the context of preceding tokens. Figure 1 illustrates this structural difference in the writing processes of LLMs and humans. We note that LLMs may also internally capture these hierarchical structures to some degree as a by-product, especially in their self-attention matrices (Xiao et al., 2021; Huber and Carenini, 2022). However, our interest lies in exploring whether there exist distinctive hierarchical structures that can aid in distinguishing their authorship. [^1] We provide a more comprehensive comparison with prior studies in Appendix A.
@@ 2.1 Rhetorical Structure Theory
RST assumes that any well-written document can be parsed into a (recursive) discourse tree. In this tree, each leaf node, called an Elementary Discourse Unit (EDU), corresponds to a phrase within a sentence, while higher-level nodes integrate these lower-level representations into more complex structures, such as phrases to sentences to paragraphs. It does so by assigning a discourse label (e.g., elaboration, contrast, cause, effect, etc.) to a relation (i.e., edge) between nodes. The linked nodes can be either “nuclei” (main ideas) or “satellites” (supporting details) depending on their relative importance in the relation. Figure 3 illustrates how a quote from Steve Jobs is segmented into seven EDUs. [^2] Full RST tree can be found in Figure 8 in Appendix B.
@@ 2.2 Conversion to Hypergraphs
Figure 3: A quote from Steve Jobs and its RST tree (Fig. 8) converted into a hypergraph format. A hexagonal node represents the “nucleus” node, while a circular one denotes the “satellite” node. Each node is labeled with a span of EDU indices that it covers. The 8-point star-shaped node is the root node of the graph, encompassing all subgraphs and EDUs.
Within the RST framework, discourse relations emerge not only through the linking of EDUs but also by encapsulating other relations. To naturally represent this process, we opt to utilize recursive hypergraphs, also referred to as “Ubergraphs” in mathematics (Joslyn and Nowak, 2017), where a discourse relation aligns with a recursive hyperedge. This decision is motivated by the potential to leverage various graph analysis techniques through the substitution of the recursive hypergraph with an isomorphic Levi graph.
In short, hypergraphs generalize traditional graphs by allowing edges (or “hyperedges”) to connect more than two nodes. The recursive hypergraphs further generalize the standard hypergraphs by allowing hyperedges to contain not only nodes but also other hyperedges. This leads to a hierarchical structure, where hyperedges at different levels of recursion represent relationships at varying levels of abstraction.
To facilitate ease of analysis, the hypergraph can be transformed into a standard traditional graph form by introducing a dummy edge that (depicted as ‘hyp.’ in Figure 3), signifying the (expanded) hyperedge relation. This enables us to run standard network analysis algorithms as well as train Graph Attention Networks (GATs) (Veličković et al., 2018), leveraging the inherent discourse structure and semantics of documents.
@@ 3.1 Discourse Motifs as Unions of Triads
With this in mind, we start by generating all possible motifs with 3 vertices, i.e., “triads.” This is a minimal form of motifs other than the single edge case, “dyad.” We select the triads to be the basis motifs as we find that when converting a recursive RST tree into a standard graph, it always forms the union of subgraphs with 3 vertices. We highlight the proof in Appendix C.
This process results in 69 non-isomorphic motifs. To create larger motifs, we pick two triads from all possible pairs, join them at all possible nodes, and only keep the non-isomorphic ones, resulting in unions of two triads. We term these configurations as “double-triads.” We conduct this process one more round with pairs of a (single-)triad and a double-triad to produce a “triple-triad.” [^4] As we are joining motifs at their nodes, the total number of nodes for double-triads and triple-triads will be smaller than 6 and 9, respectively. We note that depending on how these multiples of triads are joined together, we can form discourse motifs with varying numbers of vertices such as 4, 5, and 7. Table 1 shows the number of identified motifs for each type. Graphical examples of motifs can be found in Appendix L.
@@ 3.2 Metrics for Analyses
Once we gather potentially useful discourse motifs, we simply count the number of isomorphic subgraphs for each motif across all documents in the datasets. However, as the graph isomorphism test is (possibly) NP-intermediate, the process requires heavy engineering feats to maximize computational efficiency, such as hashing the subgraphs (Shervashidze et al., 2011) and multiprocessing the datasets. [^5] The code, as well as the processed datasets for experiments, will be released for ease of reproducibility. Afterward, we can identify distinctive motifs by examining their difference distributions based on their normalized counts (i.e., frequencies) (§3.2.1) as well as calculating our proposed metric termed “motif frequency-inverse document frequency (MF-IDF)” measure (§3.2.2). The rationale behind devising the MF-IDF scoring lies in pinpointing and prioritizing significant motifs, as counting the entire motif distribution is inefficient. An analogy can be drawn to crafting a localized tokenization scheme for LLM pre-training, wherein tokens represent motifs, and texts serve as hypergraphs in our context.
@@ 3.2.1 Motif Difference Distributions
To better capture the hierarchical nature of discourse graphs, we also calculate the average depths of each motif $m$ appearing in a graph $g$. Let $S(m,g)$ be a collection of subgraphs of a graph $g$ that are isomorphic to a motif $m$. The depth of a motif $m$ in a graph $g$ is measured by the mean position
@@ 4.1 Baseline Models
To evaluate the usefulness of discourse motifs, we integrate them into three baseline models for authorship detection (§5.1): Random Forest (RF) (Breiman, 2001), Graph Attention Network (GAT) (Veličković et al., 2018), and Longformer (LF) (Beltagy et al., 2020). Figure 4 depicts an overall process of various types of input features being handled by the different models with varying numbers of model parameters and granularity of inputs. RF is chosen to evaluate the baseline performance solely based on the discourse motif features without considering the semantic content of the texts. GAT receives the EDU-level representation of the texts as initial node embeddings, along with the discourse edge labels as edge embeddings. [^6] Texts are embedded by using Sentence-BERT (Reimers and Gurevych, 2019) model, all-MiniLM-L6-v2. The discourse motif features are then appended to the mean-pooled representation of the graph. LF is initialized with pretrained weights777allenai/longformer-base-4096, available on https://huggingface.co. and fed with the entire sequence of texts, in addition to the discourse motif features concatenated to its [CLS] token.
We follow Li et al. (2023) in selecting LF and its pretrained weights, but with three distinctions: (i) training a single classifier across all ten domains in the DeepfakeTextDetect dataset, (ii) using a reduced amount of training data ($\approx 56\%$ of the original set) to address data imbalance, and (iii) refraining from adjusting the decision boundary over ten domain-specific classifiers. Our objective is to create a general classifier, assessing discourse motifs across different model architectures, rather than optimizing for each domain.
@@ 4.2 Datasets
Existing datasets for the authorship detection task primarily contain short documents of 200 to 500 tokens, limiting the capture of long-term discourse relations; to overcome this limitation, we construct exceptionally lengthy generations based on 20 fictional stories on Project Gutenberg published in early November 2023.888https://www.gutenberg.org
@@
    Unconstrained: We provide the first human-written paragraph of a story and instruct the LLM to iteratively generate continuations of the story. [^9] Due to the limited input length of the existing discourse parser, we did not use the unconstrained generations in experiments.
@@
These settings are designed to enable the observation of long-term discourse patterns while varying the constraints applied to the original contexts. We generate content up to 8K tokens ($\approx$10 A4 pages) in an iterative manner, continuing from the previously generated texts. [^10] More detail on the construction can be found in Appx. F.
@@ 5 Results
Table 3: Overall detection F1 scores of models on the benchmark datasets.
@@
Figure 5: Difference distribution of motifs for TenPageStories under the fill-in-the-gap settings. The x-axis represents unique indices of single-triads while the y-axis shows the difference in motif frequency (scaled by 1e-3) of machine-generated and human-written texts for each motif. Discourse motifs indexed at 0 (Elaboration), 5 (Joint), and 28 (Temporal) seem to be useful in distinguishing the two groups.
Similarly, Figure 5 elucidates the distinct motif distributions within TenPageStories across various settings. We can see that the motifs indexed at 0 (Elaboration) and 5 (Joint) seem useful in detecting machine- and human-written texts, respectively. This may imply that human authors tend to construct narratives with a higher prevalence of Joint relations, indicative of more evenly branching structures, in contrast to the patterns found in texts generated by LLMs. Another noteworthy observation is the heightened occurrence of the motif indexed at 28 (Temporal) with increasing text length to generate.
@@ 5.4 Other Experiments
We conduct preliminary analyses on discourse motifs and investigate the correlation between document graph shapes across different domains in Appendices E and I, respectively.
@@ Appendix A Comparison with prior work
Appendix B An example of RST tree for the Steve Job’s quote
The RST tree in the figure is drawn with a web tool created by Zeldes (2016).
Appendix C Recursive hypergraphs to their standard traditional graph form
In this section, we introduce how recursive hypergraph representations of RST trees can be transformed into standard traditional graph forms. Suppose an RST tree consists of $m$ EDUs $\{\textrm{EDU}_{1},\cdots,\textrm{EDU}_{m}\}$ with their relations such as $(\textrm{EDU}_{1},\textrm{EDU}_{2})$ and $((\textrm{EDU}_{1},\textrm{EDU}_{2}),\textrm{EDU}_{3})$ (see EDU1, EDU2, EDU3 in Figure 8 for example). Note that in this section we ignore edge labels for better readability. A recursive hypergraph representation of the given RST in Figure 8 is a tuple $(V,E)$ where a set of vertices $V=\{\textrm{EDU}_{1},\cdots,\textrm{EDU}_{7}\}$ and a set of edges $E=\{e_{1-2}=(\textrm{EDU}_{1},\textrm{EDU}_{2}),e_{1-3}=(e_{1-2},\textrm{EDU}_{3}),\cdots,e_{1-7}=(e_{1-3},e_{4-7})\}$. [^11] We can consider bi-direction by adding a reverse order of an edge. To facilitate ease of analysis, the hypergraph can be transformed into a standard traditional graph form as follows.
@@ Appendix C Recursive hypergraphs to their standard traditional graph form
In the definition of the transformed graph, each relation in an RST tree constitutes a triangle graph. A triangle graph is a graph with three vertices where there is at least one edge between two vertices. For example, a relation $r=(x,y)$ ($x$ and $y$ are either an EDU or an edge) is transformed to the triangle graph $V^{\prime}=\{r,x,y\},E^{\prime}=\{(x,y),(x,r),(y,r)\}$. Now, we generalize this observation as follows.
@@
Proof sketch by the mathematical induction. Before we prove, assume that every finite RST can be represented by a huge, single relation in a recursive manner. For example, the RST in Figure 8 can be represented by
@@
More generally, suppose that an RST tree with less than $n$ EDUs forms the union of triangle graphs. For an RST tree $z$ containing $n$ EDUs, this RST can be divided into an RST tree $x$ with $n-k$ EDUs and an RST tree $y$ with $k$ EDUs, i.e., $z=(x,y)$. Since both transformed graphs of $x$ and $y$ are triangle unions by the assumption and a relation $(x,y)$ can be transformed into the triangle graph as discussed above, the RST $z$ can be also transformed into a new graph with triangle unions.
Appendix D Calculation of motif features

### v2

@@ Abstract
With the advent of large language models (LLM), the line between human-crafted and machine-generated texts has become increasingly blurred. This paper delves into the inquiry of identifying discernible and unique linguistic properties in texts that were written by humans, particularly uncovering the underlying discourse structures of texts beyond their surface structures. Introducing a novel methodology, we leverage hierarchical parse trees and recursive hypergraphs to unveil distinctive discourse patterns in texts produced by both LLMs and humans. Empirical findings demonstrate that, although both LLMs and humans generate distinct discourse patterns influenced by specific domains, human-written texts exhibit more structural variability, reflecting the nuanced nature of human writing in different domains. Notably, incorporating hierarchical discourse features enhances binary classifiers’ overall performance in distinguishing between human-written and machine-generated texts, even on out-of-distribution and paraphrased samples. This underscores the significance of incorporating hierarchical discourse features in the analysis of text patterns. The code and dataset are available at https://github.com/minnesotanlp/threads-of-subtlety.
@@
Zae Myung Kim1  and Kwang Hee Lee2  and Preston Zhu1  and Vipul Raheja3  and Dongyeop Kang1 University of Minnesota Twin Cities1, Kumoh National Institute of Technology2, Grammarly3 {kim01756,zhu00604,dongyeop}@umn.edu, kwanghee@kumoh.ac.kr, vipul.raheja@grammarly.com
@@ 1 Introduction
On the contrary, LLMs, in consideration, are autoregressive models that generate the next token based on the previous sequence, i.e. $P(x_{t}|x_{1}..x_{t-1})$, without explicitly modeling the hierarchical structures in the process. Figure 1 illustrates this structural difference in the writing processes of LLMs and humans. We note that LLMs may also internally capture these hierarchical structures to some degree as a by-product, especially in their self-attention matrices (Xiao et al., 2021; Huber and Carenini, 2022). However, our interest lies in exploring whether there exist distinctive hierarchical structures that can aid in distinguishing their authorship. [^1] We provide a more comprehensive comparison with prior studies in Appendix A.
@@
Figure 3: A quote from Steve Jobs and its RST tree converted into a hypergraph form. A hexagonal node represents the “nucleus” node, while a circular one denotes the “satellite” node. Each node is labeled with a span of EDU indices that it covers. The star-shaped node is the root node of the graph, encompassing all subgraphs and EDUs.
@@
  • We demonstrate that hierarchical discourse structures, as defined by RST, can be effectively modeled using recursive hypergraphs.
@@ 2.1 Rhetorical Structure Theory
RST assumes that any well-written document can be parsed into a (recursive) discourse tree. In this tree, each leaf node, called an Elementary Discourse Unit (EDU), corresponds to a phrase within a sentence, while higher-level nodes integrate these lower-level representations into more complex structures, such as phrases to sentences to paragraphs. It does so by assigning a discourse label (e.g., elaboration, contrast, cause, effect, etc.) to a relation (i.e., edge) between nodes. The linked nodes can be either “nuclei” (main ideas) or “satellites” (supporting details) depending on their relative importance in the relation. Figure 3 illustrates how a quote from Steve Jobs is segmented into seven EDUs, each marked in a distinct color, and parsed to form a discourse tree structure.
@@ 2.2 Conversion to Hypergraphs
Within the RST framework, discourse relations emerge not only through the linking of EDUs but also by encapsulating other relations. Specifically, RST relations are applied to a text “recursively” until every unit within the text becomes a constituent of an RST relation. This recursive application implies that a single EDU can form an RST relation with another individual EDU, and also become part of a relation involving a group of EDUs. For example, in Figure 3, EDU 4 and EDU 5 are linked by a “condition” relation, while EDU 6 forms a “joint” relation with the collective unit of EDU 4 and EDU 5. Additionally, since the “joint” relation encompasses the “condition” relation, RST trees inherently demonstrate recursive encapsulations.
This recursive encapsulation characteristic of RST trees naturally corresponds with recursive hypergraphs, also referred to as “Ubergraphs” in mathematics (Joslyn and Nowak, 2017), where a discourse relation aligns with a recursive hyperedge. In short, hypergraphs generalize traditional graphs by allowing edges (or “hyperedges”) to connect more than two nodes. The recursive hypergraphs further generalize the standard hypergraphs by allowing hyperedges to contain not only nodes but also other hyperedges. This leads to a hierarchical structure, where hyperedges at different levels of recursion represent relationships at varying levels of abstraction.
To facilitate ease of analysis, the hypergraph can be transformed into an isomorphic Levi graph (i.e., a standard traditional graph form) by introducing a dummy edge (depicted as a dotted line in Figure 3) that signifies the expanded hyperedge relation. By expanding the nested recursions, we are now able to run standard network analysis algorithms as well as train Graph Attention Networks (GATs) (Veličković et al., 2018), leveraging the inherent discourse structure and semantics of documents.
We note that although Yadati (2020) briefly explored the transformation of text into recursive hypergraphs, their methodology defined hyperedge relations as the relationships between entities (such as persons, places, etc.) occurring within the sentence-level organization of the text.
@@ 3.1 Discourse Motifs as Unions of Triads
With this in mind, we start by generating all possible motifs with 3 vertices, i.e., “triads.” This is a minimal form of motifs other than the single edge case, “dyad.” We select the triads to be the basis motifs as we find that when converting a recursive RST tree into a standard graph, it always forms the union of subgraphs with 3 vertices. We highlight the proof in Appendix B.
This process results in 69 non-isomorphic motifs. To create larger motifs, we pick two triads from all possible pairs, join them at all possible nodes, and only keep the non-isomorphic ones, resulting in unions of two triads. We term these configurations as “double-triads.” We conduct this process one more round with pairs of a (single-)triad and a double-triad to produce a “triple-triad.” [^3] As we are joining motifs at their nodes, the total number of nodes for double-triads and triple-triads will be smaller than 6 and 9, respectively. We note that depending on how these multiples of triads are joined together, we can form discourse motifs with varying numbers of vertices such as 4, 5, and 7. Table 1 shows the number of identified motifs for each type. Graphical examples of motifs can be found in Appendix K.
@@ 3.2 Metrics for Analyses
Once we gather potentially useful discourse motifs, we simply count the number of isomorphic subgraphs for each motif across all documents in the datasets. However, as the graph isomorphism test is (possibly) NP-intermediate, the process requires heavy engineering feats to maximize computational efficiency, such as hashing the subgraphs (Shervashidze et al., 2011) and multiprocessing the datasets. [^4] The code and the processed datasets for the experiments can be found at https://github.com/minnesotanlp/threads-of-subtlety. Afterward, we can identify distinctive motifs by examining their difference distributions based on their normalized counts (i.e., frequencies) (§3.2.1) as well as calculating our proposed metric termed “motif frequency-inverse document frequency (MF-IDF)” measure (§3.2.2). The rationale behind devising the MF-IDF scoring lies in pinpointing and prioritizing significant motifs, as counting the entire motif distribution is inefficient. An analogy can be drawn to crafting a localized tokenization scheme for LLM pre-training, wherein tokens represent motifs, and texts serve as hypergraphs in our context.
@@ 3.2.1 Motif Difference Distributions
To better capture the hierarchical nature of discourse graphs, we also calculate the average depths of each motif $m$ appearing in a graph $g$. Let $S(m,g)$ be a collection of subgraphs of a graph $g$ that are isomorphic to a motif $m$. Then, the depth of a motif $m$ in a graph $g$ is measured by the mean position
@@ 4.1 Baseline Models
To evaluate the usefulness of discourse motifs, we integrate them into three baseline models for authorship detection (§5.1): Random Forest (RF) (Breiman, 2001), Graph Attention Network (GAT) (Veličković et al., 2018), and Longformer (LF) (Beltagy et al., 2020). Figure 4 depicts an overall process of various types of input features being handled by the different models with varying numbers of model parameters and granularity of inputs.
RF is chosen to evaluate the baseline performance solely based on the discourse motif features (defined in §3.2.1) without considering the semantic content of the texts.
GAT receives the EDU-level representation of the texts as initial node embeddings, along with the discourse edge labels as edge embeddings. The texts for both EDUs and discourse labels are embedded by using Sentence-BERT (Reimers and Gurevych, 2019) model, all-MiniLM-L6-v2. The discourse motif features are then appended to the mean-pooled representation of the graph.
LF is initialized with pretrained weights555allenai/longformer-base-4096, available on https://huggingface.co. and fed with the entire sequence of texts, in addition to the discourse motif features concatenated to its [CLS] token. Specifically, we align our use of LF with the methodology described by Li et al. (2023), which details a state-of-the-art detection model. However, we introduce three key distinctions: (i) we train a single classifier across all ten domains of the DeepfakeTextDetect dataset, (ii) we utilize approximately 56% of the original dataset to mitigate data imbalance issues, and (iii) we avoid adjusting the decision boundary across ten domain-specific classifiers. Our goal is to develop a general classifier capable of evaluating discourse motifs across various model architectures, rather than tailoring optimizations for individual domains.
We note that our approach can be seamlessly integrated into other baselines by concatenating discourse motif features with input representations. Additionally, it can enhance other detection strategies by providing extra statistical features for zero-shot detectors or by aiding in the selection of more challenging and deceptive samples for adversarial learning setups.
@@ 4.2 Datasets
Existing datasets for the authorship detection task primarily contain short documents of 200 to 500 tokens, limiting the capture of long-term discourse relations; to overcome this limitation, we construct exceptionally lengthy generations based on 20 fictional stories on Project Gutenberg published in early November 2023.666https://www.gutenberg.org
@@
    Unconstrained: We provide the first human-written paragraph of a story and instruct the LLM to iteratively generate continuations of the story. However, due to the limited input length of the existing discourse parser, we did not use the unconstrained generations in experiments.
@@
These settings are designed to enable the observation of long-term discourse patterns while varying the constraints applied to the original contexts. We generate content up to 8K tokens ($\approx$10 A4 pages) in an iterative manner, continuing from the previously generated texts. We provide more details on the dataset construction in Appendix E.
@@ 5 Results
Table 3: Overall detection F1 scores of models on the benchmark datasets. “RF (All Motifs)” denotes Random Forest models with all of the found motifs (3,055 of them) considered as inputs. “RF (Motifs)” indicates Random Forest models where only the MF-IDF-selected motifs (207 of them) are taken as inputs.
@@
Figure 5: Difference distribution of motifs for TenPageStories under the fill-in-the-gap settings. The x-axis represents unique indices of single-triads while the y-axis shows the difference in motif frequency (scaled by 1e-3) of machine-generated and human-written texts for each motif. Discourse motifs indexed at 0 (Elaboration), 5 (Joint), 7 (Joint), and 28 (Temporal) seem to be useful in distinguishing the two groups.
Similarly, Figure 5 elucidates the distinct motif distributions within TenPageStories across various settings. It shows that as the generation length increases, motifs indexed at 0 (Elaboration) and 28 (Temporal) become increasingly prevalent in machine-generated texts. Conversely, motifs indexed at 7 (Joint) and 5 (Joint) appear more frequently in texts authored by humans. This may imply that human authors tend to construct narratives with a higher prevalence of Joint relations, indicative of more evenly branching structures, in contrast to the patterns found in texts generated by LLMs.
@@ 5.4 Other Experiments
We conduct preliminary analyses on discourse motifs and investigate the correlation between document graph shapes across different domains in Appendices D and H, respectively.
@@ 8 Ethical Statement
Acknowledgements
This research was primarily funded by a generous research gift from Grammarly. Kwang Hee Lee received funding from the National Research Foundation of Korea (NRF), which is supported by the Korean government (MSIT), under Grant No. RS-2024-00345567. We are grateful to the Minnesota NLP group members for their valuable feedback and constructive comments on our initial draft. We also acknowledge Jonghwan Hyeon and Chae-Gyun Lim from KAIST for their assistance with the design of the figures.
@@ Appendix A Comparison with prior work
Appendix B Recursive hypergraphs to their standard traditional graph form
In this section, we introduce how recursive hypergraph representations of RST trees can be transformed into standard traditional graph forms. Suppose an RST tree consists of $m$ EDUs $\{\textrm{EDU}_{1},\cdots,\textrm{EDU}_{m}\}$ with their relations such as $(\textrm{EDU}_{1},\textrm{EDU}_{2})$ and $((\textrm{EDU}_{1},\textrm{EDU}_{2}),\textrm{EDU}_{3})$ (see EDU1, EDU2, EDU3 in Figure 8 for example). Note that in this section, we ignore edge labels for better readability. A recursive hypergraph representation of the given RST in Figure 8 is a tuple $(V,E)$ where a set of vertices $V=\{\textrm{EDU}_{1},\cdots,\textrm{EDU}_{7}\}$ and a set of edges $E=\{e_{1-2}=(\textrm{EDU}_{1},\textrm{EDU}_{2}),e_{1-3}=(e_{1-2},\textrm{EDU}_{3}),\cdots,e_{1-7}=(e_{1-3},e_{4-7})\}$. [^7] We can consider bi-direction by adding a reverse order of an edge. To facilitate ease of analysis, the hypergraph can be transformed into a standard traditional graph form as follows.
@@ Appendix B Recursive hypergraphs to their standard traditional graph form
In the definition of the transformed graph, each relation in an RST tree constitutes a triangle graph. A triangle graph is a graph with three vertices where there is at least one edge between two vertices. For example, a relation $r=(x,y)$ ($x$ and $y$ are either an EDU or an edge) is transformed to the triangle graph, $V^{\prime}=\{r,x,y\},E^{\prime}=\{(x,y),(x,r),(y,r)\}$. Now, we generalize this observation as follows.
@@
Proof sketch by induction. Before we prove, assume that every finite RST can be represented by a huge, single relation in a recursive manner. For example, the RST in Figure 8 can be represented by
@@
More generally, suppose that an RST tree with less than $n$ EDUs forms the union of triangle graphs. For an RST tree $z$ containing $n$ EDUs, this RST can be divided into an RST tree $x$ with $n-k$ EDUs and an RST tree $y$ with $k$ EDUs, i.e., $z=(x,y)$. Since both transformed graphs of $x$ and $y$ are unions of triangle graphs by the assumption, and a relation $(x,y)$ can be transformed into the triangle graph as discussed above, the RST $z$ can be also transformed into a new graph with unions of triangle graphs.
Appendix C Calculation of motif features